OpenStudy (anonymous):

Solve for x: 81^x = 27^-x

jimthompson5910 (jim_thompson5910):

81^x = 27^(-x) (3^4)^x = (3^3)^(-x) 3^(4x) = 3^(-3x) since the bases are equal, the exponents are equal, so 4x = -3x 4x+3x = 0 7x = 0 x = 0/7 x = 0

OpenStudy (anonymous):

OMG Ty !

jimthompson5910 (jim_thompson5910):

yw